OpenStudy (anonymous):

What is the magnitude and direction of Vector GH if G(2, -2) and H(7, 6)?

OpenStudy (anonymous):

I have the multiple choice answers if anyone needs it, I just really want to understand how to solve it.

OpenStudy (waleed_imtiaz):

TO find the magnitude,,, U have to apply the formula..... |GH| = \[\sqrt{(x2-x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (waleed_imtiaz):

Can U solve this ?

OpenStudy (anonymous):

Yeah, that's distance formula. square root of 89

OpenStudy (anonymous):

so, the direction. How do I find it?

OpenStudy (ghazi):

write vector GO= 2i-2j and another vector HO= 7i-6j .....now your final vector would be = GO-HO= (7i-6j)-(2i-2j)....after that find the mod of resultant vector that is = \[\sqrt{(a)^2+(b)^2}\] where a and b is coefficient of i and j..you'll get both direction and magnitude

OpenStudy (anonymous):

Wait, what is GO and i, j?

OpenStudy (phi):